OpenStudy (anonymous):

F(t)= 98 + 4/(t+1) find F'(t), F(3), F'(3) F' is the prime

OpenStudy (anonymous):

F'(t)=-4/(t+1)^2==>derivative of a quotient plugging in F(3)=98+4/(3+1)=99 plugging in F'(3)=-1/4